In infinite dilutions,the equivalent conductances of Ba2+ and Cl- are 127 and 76 respectively. The equivalent conductivity of BaCl2 at indefinite dilution is 
(a) 101.5  (b) 139.5  (ć) 203.5  (d) 279.5

Grade:12th pass

1 Answers

Arun
5 years ago
The conductance is Ba2+ + 2Cl-
= 127+ 2(76)

= 279 /2 = 139.5 /ohm cm2
 
 
Regards
Arun (askIITiaaskIITians forforum expert)